OpenStudy (anonymous):

Is the square root of x-3 a one to one function? I believe it is, but I'm unsure.

OpenStudy (aum):

Yes, it is. It passes the horizontal line test.

OpenStudy (aum):



OpenStudy (anonymous):

Thank you!

OpenStudy (zzr0ck3r):

yes suppose \(\sqrt{x-3}=\sqrt{y-3}\implies x-3=y-3\implies x=y\)

OpenStudy (aum):

No two values of x gives the same y value. Therefore, it is a one-to-one function. You are welcome.